Operations Report – May 14th

Crew 194 Operations Report 14052018

SOL: 1

Name of person filing report: Dana

Non-nominal systems: Power/ Generator

Notes on non-nominal system:

functional but does not charge batteries

Solar—insufficient to power hab but does appear connected to batteries

Diesel – 75%

Gasoline (5 Gallon containers for ATV) – 15 gallons

Water (trailer) – 920 gallons

Water (static) – 55 gallons

Trailer to Static Pump used – no

Water (loft) – Static to Loft Pump used – yes

Water Meter: nominal

ATV’s Used: 4

Oil Checked: yes

ATV # Fuel Used Gals: 7 gallons

Tires Status: good

# Hours Used Day: 2

Phobos used: yes

Hours: 2

Beginning charge: 100%

Ending charge: 95%

Currently charging: yes

Deimos used: No

Hours:

Beginning charge:

Ending charge:

Currently charging:

Sojourner rover used: ASSIGNED TO DIRECTOR

Hours:

Beginning charge:

Ending charge:

Currently charging:

Spirit rover used: No

Hours:

Beginning charge:

Ending charge:

Currently charging:

Opportunity rover used: No

Hours:

Beginning charge:

Ending charge:

Currently charging:

Curiosity rover used: No

Hours:

Beginning charge:

Ending charge:

Currently charging:

HabCar used? No

Notes and Comments:

Summary of operations activities:

Summary of GreenHab activities: harvested Tomatoes and basil

Summary of ScienceDome activities: worked on construction of simulated Rocket

Summary of healthy and safety issues: none

Questions and Concerns to Mission Support: Power system remains off nominal. We are able to maintain electricity by using the diesel generator but the governor electrical system continues to not function.

Operations Report – May 5th

Crew 193 Operations Report 05May2018

SOL over

Name of person filing report: David Attig

Non-nominal systems: Two piece suits still need straps adjusted for neck ring. Generator/inverter/charger. One piece suit #9
Notes on non-nominal systems: The generator still appears to not be charging the batteries at night (only powering the HAB). Shannon is having a look, this should not affect our operations (particularly with the days being so long now).

Generator (hours run): 8:45 pm to 7:30 am.

Solar— SOC 5 % (Before generator is run last night)

Diesel – 60 %

Propane – 60 psi.

ATV Fuel (Ethanol free gasoline) – 0 gallons.

Water (trailer) – 500 gallons.

Water (static) – 500 gallons

Trailer to Static Pump used – no

Water (loft) – Static to Loft Pump used – yes

Water Meter: 134635.2

Toilet tank emptied: no
Date last emptied: May 2

ATV’s (Honda, 300, 350.1, 350.2, 350.3): No

ATV Oil Checked?: No

ATV Oil Added?: No

ATV Fuel Used: 0 Gallons

# Hours the ATVs were Used today: 0

Notes on ATVs: 350.2 is out of service, due to a previous issue. Honda has a hole in the seat. 300 would not start, there is a grinding or clicking sound (dead battery?)

Deimos rover used: No
Hours: 125.6
Beginning charge: 100%
Ending charge: N/A
Currently charging: No

Sojourner rover used: ASSIGNED TO DIRECTOR
Hours: Unknown
Beginning charge: N/A
Ending charge: N/A
Currently charging: N/A

Spirit rover used: Yes
Hours: 36.0
Beginning charge: 100 %
Ending charge: 99 %
Currently charging: Yes

Opportunity rover used: Yes
Hours: 31.6
Beginning charge: 100 %
Ending charge: 98 %
Currently charging: Yes

Curiosity rover used: Yes
Hours: 29.9
Beginning charge: 100 %
Ending charge: 98 %
Currently charging: No

Hab Car used?: Yes
Where was Hab Car used?: gas station
Why was Hab Car used?: Water replenishment

General notes and comments: none

Summary of internet: We’re out, same as normal.
Summary of suits and radios: Two piece suits still need strap maintenance

Summary of Hab operations: Super nominal.

Summary of GreenHab operations: Water saturated

Summary of ScienceDome operations: Concluded

Summary of RAM operations: cleaned

Summary of health and safety issues: None to report

Questions, concerns and requests to Mission Support: Sol over

Operations Report – May 04th

SOL: 12

Name of person filing report: David Attig

Non-nominal systems: Two piece suits still need straps adjusted for neck ring. Generator/inverter/charger. One piece suit #10.

Notes on non-nominal systems: The generator still appears to not be charging the batteries at night (only powering the HAB). Shannon is having a look, this should not affect our operations (particularly with the days being so long now).

Generator (hours run): 8:45 pm to 7:25 am.

Solar— SOC 60 % (Before generator is run at night)

Diesel – 60 %

Propane – 60 psi.

ATV Fuel (Ethanol free gasoline) – 0 gallons.

Water (trailer) – 30 gallons.

Water (static) – 285 gallons

Trailer to Static Pump used – no

Water (loft) – Static to Loft Pump used – yes

Water Meter: 134635.2

Toilet tank emptied: no
Date last emptied: May 2

ATV’s (Honda, 300, 350.1, 350.2, 350.3): Honda, 350.1, 350.3

ATV Oil Checked?: No

ATV Oil Added?: No

ATV Fuel Used: 0 Gallons

# Hours the ATVs were Used today: 2.5

Notes on ATVs: 350.2 is out of service, due to a previous issue. Honda has a hole in the seat. 300 would not start, there is a grinding or clicking sound, as if teeth are broken off of the starter or flywheel.

Deimos rover used: Yes
Hours: 125.6
Beginning charge: 100%
Ending charge: 73 %
Currently charging: Yes

Sojourner rover used: ASSIGNED TO DIRECTOR
Hours: Unknown
Beginning charge: N/A
Ending charge: N/A
Currently charging: N/A

Spirit rover used: Yes
Hours: 35.8
Beginning charge: 100 %
Ending charge: 86 %
Currently charging: Yes

Opportunity rover used: No
Hours: 34.2
Beginning charge: 100 %
Ending charge: N/A %
Currently charging: No

Curiosity rover used: No
Hours: 29.8
Beginning charge: 100 %
Ending charge: N/A %
Currently charging: No

Hab Car used?: No
Where was Hab Car used?: N/A
Why was Hab Car used?: N/A

General notes and comments: Supported two EVAs today.

Summary of internet: We’re out, same as normal.

Summary of suits and radios: Still work to do on two piece suits (straps). Suit # 7 failed during the final EVA today; fans got stuck in shroud and released magic smoke.

Summary of Hab operations: Super nominal.

Summary of GreenHab operations: Watered twice.

Summary of ScienceDome operations: Dishwasher experiment concluding.

Summary of RAM operations: Outreach video created in RAM today.

Summary of health and safety issues: None to report

Questions, concerns and requests to Mission Support: Our garbage is ready to be picked up (both burnable and non burnable). We would also like more ATV fuel (ethanol free gasoline).

Operations Report – 03 May 2018

SOL: 11

Name of person filing report: David Attig

Non-nominal systems: Two piece suits still need straps adjusted for neck ring. Generator/inverter/charger.
Notes on non-nominal systems: The generator still appears to not be charging the batteries at night (only powering the HAB). Shannon is having a look, this should not affect our operations (particularly with the days being so long now).

Generator (hours run): 9:00 pm to 9:15 am.

Solar— SOC 19 % (Before generator is run at night)

Diesel – 65 %

Propane – 60 psi.

ATV Fuel (Ethanol free gasoline) – 0 gallons.

Water (trailer) – 30 gallons.

Water (static) – 350 gallons

Trailer to Static Pump used – yes

Water (loft) – Static to Loft Pump used – yes

Water Meter: 134569.7

Toilet tank emptied: no
Date last emptied: May 2

ATV’s (Honda, 300, 350.1, 350.2, 350.3): 350.1

ATV Oil Checked?: No

ATV Oil Added?: No

ATV Fuel Used: 0 Gallons

# Hours the ATVs were Used today: 0.1

Notes on ATVs: 350.2 is out of service, due to a previous issue. Honda has a hole in the seat.

Deimos rover used: No
Hours: 124.8
Beginning charge: 100%
Ending charge: N/A %
Currently charging: No

Sojourner rover used: ASSIGNED TO DIRECTOR
Hours: Unknown
Beginning charge: N/A
Ending charge: N/A
Currently charging: N/A

Spirit rover used: No
Hours: 35.0
Beginning charge: 100 %
Ending charge: N/A %
Currently charging: No

Opportunity rover used: No
Hours: 34.2
Beginning charge: 100 %
Ending charge: N/A %
Currently charging: Yes

Curiosity rover used: No
Hours: 29.8
Beginning charge: 100 %
Ending charge: N/A %
Currently charging: Yes

Hab Car used?: No
Where was Hab Car used?: N/A
Why was Hab Car used?: N/A

General notes and comments: Had visitors to HAB today, spent most time giving a tour

Summary of internet: We’re out, same as normal.
Summary of suits and radios: One piece suit #10 is repaired. Still work to do on two piece suits.

Summary of Hab operations: Muddily nominal

Summary of GreenHab operations: Nominal; watered twice

Summary of ScienceDome operations: Nominal

Summary of RAM operations: Children ate lunch inside RAM, CO secured after their visit.

Summary of health and safety issues: None to report

Questions, concerns and requests to Mission Support: Nothing yet.

Operations Report – May 2nd

Crew 193 Operations Report 02May2018

SOL: 10

Name of person filing report: David Attig

Non-nominal systems: Two piece suits still need straps adjusted for neck ring. Generator/inverter/charger. One piece suit #10.

Notes on non-nominal systems: The generator still appears to not be charging the batteries at night (only powering the HAB). Shannon is having a look, this should not affect our operations (particularly with the days being so long now).

Generator (hours run): 7:00 pm to 7:55 am.

Solar— SOC 60 % (Before generator is run at night)

Diesel – 70 %

Propane – 60 psi.

ATV Fuel (Ethanol free gasoline) – 0 gallons.

Water (trailer) – 220 gallons.

Water (static) – 290 gallons

Trailer to Static Pump used – no

Water (loft) – Static to Loft Pump used – yes

Water Meter: 134505.5

Toilet tank emptied: yes
Date last emptied: May 2

ATV’s (Honda, 300, 350.1, 350.2, 350.3): No

ATV Oil Checked?: No

ATV Oil Added?: No

ATV Fuel Used: 0 Gallons

# Hours the ATVs were Used today: 0

Notes on ATVs: 350.2 is out of service, due to a previous issue. Honda has a hole in the seat.

Deimos rover used: No
Hours: 124.8
Beginning charge: 100%
Ending charge: N/A %
Currently charging: No

Sojourner rover used: ASSIGNED TO DIRECTOR
Hours: Unknown
Beginning charge: N/A
Ending charge: N/A
Currently charging: N/A

Spirit rover used: No
Hours: 35.0
Beginning charge: 100 %
Ending charge: N/A %
Currently charging: No

Opportunity rover used: No
Hours: 34.2
Beginning charge: 100 %
Ending charge: N/A %
Currently charging: Yes

Curiosity rover used: No
Hours: 29.8
Beginning charge: 100 %
Ending charge: N/A %
Currently charging: Yes

Hab Car used?: No
Where was Hab Car used?: N/A
Why was Hab Car used?: N/A

General notes and comments: Muddy

Summary of internet: We’re out, same as normal.

Summary of suits and radios: One piece suit #10 is repaired. Still work to do on two piece suits. One suit charger repaired, repair of other suit chargers still in progress.

Summary of Hab operations: Muddily nominal

Summary of GreenHab operations: Muddy

Summary of ScienceDome operations: Muddy

Summary of RAM operations: None today, Muddy

Summary of health and safety issues: Muddy

Questions, concerns and requests to Mission Support: Less mud. And an escalator.

Operations Report – May 1st

Crew 193 Operations Report 01May2018

SOL: 9

Name of person filing report: David Attig

Non-nominal systems: Two piece suits still need straps adjusted for neck ring. Generator/inverter/charger. One piece suit #10.
Notes on non-nominal systems: The generator still appears to not be charging the batteries at night (only powering the HAB). Shannon is having a look, this should not affect our operations (particularly with the days being so long now).

Generator (hours run): 7:00 pm to 7:55 am.

Solar— SOC 60 % (Before generator is run at night)

Diesel – 70 %

Propane – 60 psi.

ATV Fuel (Ethanol free gasoline) – 0 gallons.

Water (trailer) – 220 gallons.

Water (static) – 345 gallons

Trailer to Static Pump used – no

Water (loft) – Static to Loft Pump used – yes

Water Meter: 134437.8

Toilet tank emptied: no
Date last emptied: April 29

ATV’s (Honda, 300, 350.1, 350.2, 350.3): Honda, 300, 350.1, 350.3

ATV Oil Checked?: No

ATV Oil Added?: No

ATV Fuel Used: 0.5 Gallons

# Hours the ATVs were Used today: 3

Notes on ATVs: 350.2 is out of service, due to a previous issue. Honda has a hole in the seat.

Deimos rover used: No
Hours: 124.8
Beginning charge: 100%
Ending charge: N/A %
Currently charging: No

Sojourner rover used: ASSIGNED TO DIRECTOR
Hours: Unknown
Beginning charge: N/A
Ending charge: N/A
Currently charging: N/A

Spirit rover used: No
Hours: 35.0
Beginning charge: 100 %
Ending charge: N/A %
Currently charging: No

Opportunity rover used: No
Hours: 34.2
Beginning charge: 100 %
Ending charge: N/A %
Currently charging: Yes

Curiosity rover used: No
Hours: 29.8
Beginning charge: 100 %
Ending charge: N/A %
Currently charging: Yes

Hab Car used?: No
Where was Hab Car used?: N/A
Why was Hab Car used?: N/A

General notes and comments: Supported two EVAs today.

Summary of internet: We’re out, same as normal.
Summary of suits and radios: Still work to do on two piece suits. Repair of suit chargers still in progress. Other than #10, one piece suits nominal.

Summary of Hab operations: Rear airlock door was pulled on while latched (we accidentally locked a crew-member out). While pulling, one screw came loose; it has been replaced by a screw with more length and girth.

Summary of GreenHab operations: Watered twice.

Summary of ScienceDome operations: Dishwasher experiment proceeding.

Summary of RAM operations: None today

Summary of health and safety issues: None to report

Questions, concerns and requests to Mission Support: Our garbage is ready to be picked up (both burnable and non burnable). There is also cardboard from sawhorse and wagon boxes. We would also like more ATV fuel (ethanol free gasoline).

Operations Report – April 30th

Crew 193 Operations Report 30Apr2018

SOL: 8

Name of person filing report: David Attig

Non-nominal systems: Two piece suits still need straps adjusted for neck ring. Generator/inverter/charger. One piece suit #10. Soldering Irons

Notes on non-nominal systems: The generator still appears to not be charging the batteries at night (only powering the HAB). Shannon is having a look, this should not affect our operations (particularly with the days being so long now). At 4:45 PM today, the SoC was 73%, which is lower than it should be, given how sunny it is today; there are some storm clouds to the southwest, which could be blocking the sun, we need to keep an eye on SoC.
I will repair charging plugs once we have a soldering iron.

Generator (hours run): 10:45 pm to 8:10 am.

Solar— SOC 59 % (Before generator is run at night)

Diesel – 70 %

Propane – 60 psi.

ATV Fuel (Ethanol free gasoline) – 0.5 gallons.

Water (trailer) – 220 gallons.

Water (static) – 400 gallons

Trailer to Static Pump used – no

Water (loft) – Static to Loft Pump used – yes

Water Meter: 134370.9

Toilet tank emptied: no
Date last emptied: April 29

ATV’s (Honda, 300, 350.1, 350.2, 350.3): Honda, 300, 350.1, 350.3

ATV Oil Checked?: No

ATV Oil Added?: No

ATV Fuel Used: 2 Gallons

# Hours the ATVs were Used today: 3

Notes on ATVs: 350.2 is out of service, due to a previous issue. Honda has a hole in the seat.

Deimos rover used: No
Hours: 124.8
Beginning charge: 100%
Ending charge: N/A %
Currently charging: No

Sojourner rover used: ASSIGNED TO DIRECTOR
Hours: Unknown
Beginning charge: N/A
Ending charge: N/A
Currently charging: N/A

Spirit rover used: No
Hours: 35.0
Beginning charge: 100 %
Ending charge: N/A %
Currently charging: No

Opportunity rover used: No
Hours: 34.2
Beginning charge: 100 %
Ending charge: N/A %
Currently charging: Yes

Curiosity rover used: No
Hours: 29.8
Beginning charge: 100 %
Ending charge: N/A %
Currently charging: Yes

Hab Car used?: No
Where was Hab Car used?: N/A
Why was Hab Car used?: N/A

General notes and comments: Supported two EVAs today.

Summary of internet: We’re out, same as normal.

Summary of suits and radios: Still work to do on two piece suits. Repair of suit chargers still in progress. Other than #10, one piece suits nominal.

Summary of Hab operations: Super nominal.

Summary of GreenHab operations: Watered twice.

Summary of ScienceDome operations: Dishwasher experiment proceeding.

Summary of RAM operations: A charging connector was moved from RAM for space suit repair (once soldering iron arrives)

Summary of health and safety issues: None to report

Questions, concerns and requests to Mission Support: We would like some black straps used for connecting two piece suit neck rings to backpacks. I did not see any in the RAM.

Operations Report – April 29th

Crew 193 Operations Report 29Apr2018

SOL: 7

Name of person filing report: David Attig

Non-nominal systems: Two piece suits still need straps adjusted for neck ring. Generator/inverter/charger. One piece suit #10. Soldering Irons
Notes on non-nominal systems: The generator still appears to not be charging the batteries at night (only powering the HAB). Shannon is having a look, this should not affect our operations (particularly with the days being so long now).
I will repair charging plugs once we have a soldering iron.
One piece suit #10’s fans stopped when I bent over while wearing the suit. The loose connection was determined to be at the positive terminal of the battery; upon inspection, there was a cold solder joint on a terminal connector. While removing tape to access battery terminals, I broke the small wires on negative terminal. Suit #10 is out of service until we have a working soldering iron (should be on the way).

Generator (hours run): 10:45 pm to 7:35 am.

Solar— SOC 57 % (Before generator is run at night)

Diesel – 70 %

Propane – 60 psi.

ATV Fuel (Ethanol free gasoline) – 2.5 gallons.

Water (trailer) – 220 gallons.

Water (static) – 455 gallons

Trailer to Static Pump used – no

Water (loft) – Static to Loft Pump used – yes

Water Meter: 134315.1

Toilet tank emptied: yes
Date last emptied: April 29

ATV’s (Honda, 300, 350.1, 350.2, 350.3): Honda, 300, 350.3

ATV Oil Checked?: Yes

ATV Oil Added?: No

ATV Fuel Used: 2.5 Gallons

# Hours the ATVs were Used today: 2

Notes on ATVs: 350.2 is out of service, due to a previous issue. Honda oil is within range, but on low end.

Deimos rover used: No
Hours: 124.8
Beginning charge: 100%
Ending charge: N/A %
Currently charging: No

Sojourner rover used: ASSIGNED TO DIRECTOR
Hours: Unknown
Beginning charge: N/A
Ending charge: N/A
Currently charging: N/A

Spirit rover used: No
Hours: 35.0
Beginning charge: 100 %
Ending charge: N/A %
Currently charging: No

Opportunity rover used: No
Hours: 34.2
Beginning charge: 100 %
Ending charge: N/A %
Currently charging: Yes

Curiosity rover used: No
Hours: 29.8
Beginning charge: 100 %
Ending charge: N/A %
Currently charging: Yes

Hab Car used?: No
Where was Hab Car used?: N/A
Why was Hab Car used?: N/A

General notes and comments: Supported two EVAs today.

Summary of internet: We’re out, same as normal.
Summary of suits and radios: Still work to do on two piece suits. Repair of suit chargers still in progress. Other than #10, one piece suits nominal.

Summary of Hab operations: Super nominal.

Summary of GreenHab operations: Watered twice.

Summary of ScienceDome operations: Dishwasher experiment proceeding, more samples taken today.

Summary of RAM operations: Sawhorses moved in from RAM.

Summary of health and safety issues: None to report

Questions, concerns and requests to Mission Support: Would like some new crimp-on connectors for suit battery terminals. For example: https://www.mcmaster.com/#7060K38. An assortment would probably be nice for future missions.

Operations Report – April 28th

Crew 193 Operations Report 28Apr2018

SOL: 6

Name of person filing report: David Attig

Non-nominal systems: Generator – sim suit chargers

Notes on non-nominal systems: The generator still appears to not be charging the batteries at night (only powering the HAB). Shannon is having a look, this should not affect our operations (particularly with the days being so long now).

Generator (hours run): 11:00 pm to 8:30 am.

Solar— SOC 70 % (Before generator is run at night)

Diesel – 70 %

Propane – 60 psi.

ATV Fuel (Ethanol free gasoline) – 5 gallons.

Water (trailer) – 220 gallons.

Water (static) – 500 gallons

Trailer to Static Pump used – yes

Water (loft) – Static to Loft Pump used – yes

Water Meter: 134250.5

Toilet tank emptied: no
Date last emptied: April 26

ATV’s (Honda, 300, 350.1, 350.2, 350.3): No

ATV Oil Checked?: No

ATV Oil Added?: No

ATV Fuel Used: 0 Gallons

# Hours the ATVs were Used today: 0

Notes on ATVs: 350.2 is out of service, due to a previous issue.

Deimos rover used: No
Hours: 124.8
Beginning charge: 100%
Ending charge: N/A %
Currently charging: No

Sojourner rover used: ASSIGNED TO DIRECTOR
Hours: Unknown
Beginning charge: N/A
Ending charge: N/A
Currently charging: N/A

Spirit rover used: No
Hours: 35.0
Beginning charge: 100 %
Ending charge: N/A %
Currently charging: No

Opportunity rover used: No
Hours: 34.2
Beginning charge: 100 %
Ending charge: N/A %
Currently charging: Yes

Curiosity rover used: No
Hours: 29.8
Beginning charge: 100 %
Ending charge: N/A %
Currently charging: Yes

Hab Car used?: No
Where was Hab Car used?: N/A
Why was Hab Car used?: N/A

General notes and comments: Outreach videos were made today, water transferred from storage tank to static tank. Red wagon assembled for greenhab.

Summary of internet: We’re out, same as normal.

Summary of suits and radios: One piece suits nominal, still work to do on two piece suits. Repair of suit chargers still in progress.

Summary of Hab operations: Transferred water from storage tank to static tank.

Summary of GreenHab operations: Watered twice, plants returning to health. Assembled red wagon placed inside.

Summary of ScienceDome operations: Dishwasher experiment proceeding, more samples taken today.

Summary of RAM operations: Not used today.

Summary of health and safety issues: One crewmember scraped their hand when going down the stairs (actually fell down the last one).

Questions, concerns and requests to Mission Support: We would like an ice cube tray for the freezer.

Operations Report – April 27th

SOL: 4

Name of person filing report: David Attig

Non-nominal systems: A couple sim suit chargers, and soldering iron.

Notes on non-nominal systems: A sim-suit charging cable broke yesterday, due to having cold solder joints. I attempted to repair, but the soldering Iron broke after heating briefly.

Generator (hours run): 9:30 pm to 9:00 am.

Solar— SOC 22 % (Before generator is run at night)

Diesel – 70 %

Propane – 60 psi.

ATV Fuel (Ethanol free gasoline) – 5 gallons.

Water (trailer) – 550 gallons.

Water (static) – 250 gallons

Trailer to Static Pump used – no

Water (loft) – Static to Loft Pump used – yes

Water Meter: 134192.9

Toilet tank emptied: no
Date last emptied: April 26

ATV’s (Honda, 300, 350.1, 350.2, 350.3): Yes, 300

ATV Oil Checked?: No

ATV Oil Added?: No

ATV Fuel Used: 0.01 Gallons

# Hours the ATVs were Used today: 0.1

Notes on ATVs: 350.2 is out of service, due to a previous issue.

Deimos rover used: No
Hours: 124.8
Beginning charge: 100%
Ending charge: N/A %
Currently charging: No

Sojourner rover used: ASSIGNED TO DIRECTOR
Hours: Unknown
Beginning charge: N/A
Ending charge: N/A
Currently charging: N/A

Spirit rover used: No
Hours: 35.0
Beginning charge: 100 %
Ending charge: N/A %
Currently charging: No

Opportunity rover used: Yes
Hours: 34.2
Beginning charge: 100 %
Ending charge: 96 %
Currently charging: Yes

Curiosity rover used: No
Hours: 29.8
Beginning charge: 100 %
Ending charge: N/A %
Currently charging: Yes

Hab Car used?: No
Where was Hab Car used?: N/A
Why was Hab Car used?: N/A

General notes and comments: Eric made progress troubleshooting the beartooth devices by communicating with the manufacturer. The issue has not been resolved yet. Last night we waited too long to turn on the generator, SoC had dropped to 22%. This morning, it was only 26% after the generator running all night. It appears that the generator is not charging the batteries at all. SoC returned to 100% after a few hours of full sunlight. Mission Control (Shannon) is investigating further.

Summary of internet: We’re out, same as normal.
Summary of suits and radios: One piece suits nominal, still work to do on two piece suits. I changed some settings on radios we use during EVA (whisper function) to make the microphone more sensitive. Recieved new radios, opened one set. They are ready to use on EVAs. Demoted some older EVA radios to science dome, retired one radio from science dome and one from the green hab (both radios have issues).

Summary of Hab operations: Started sorting tools in toolbox, and radios/chargers. Went out and checked water tank, confirmed we are half way through it.

Summary of GreenHab operations: Heavily watered (8+ gallons) in morning at request of mission control.

Summary of ScienceDome operations: Dishwasher experiment proceeding, more samples taken today.

Summary of RAM operations: New radios retrieved from RAM airlock.

Summary of health and safety issues: During a local EVA to practice medical emergencies, we had one crewmember suffer from heat exhaustion. EVA was cut short, crewmember is feeling better.

Questions, concerns and requests to Mission Support: We would like some small cardboard boxes for organizing the HAB. They would also be good for moving material and supplies during EVAs (such as to/from RAM).
We would like some hang-tags or white cloth (gaffing?) tape to mark things as we organize the HAB.
Also, we would like some small paintbrushes (for detail work on sign), wet-wipes, and paper towels.